OpenStudy (anonymous):

solve for x: (3x/y)=(6g/b)

OpenStudy (saifoo.khan):

\[\frac{3x}{y}=\frac{6g}{b}\] cross multiply. \[x = \frac{6gy}{3b}\]

OpenStudy (anonymous):

ok thats what i thought thanks

OpenStudy (saifoo.khan):

wwelks.